# arXiv:2603.00340v1[cs.LG] 27 Feb 2026

### v1

@@ 1 Introduction
The study of human mobility patterns—how individuals move across space—has become an important topic across geography, transportation science, public health, and climate change science [19; 50; 67; 51; 8; 38; 5; 75; 23; 58]. One key aspect of human mobility is transportation mode choice, the accurate estimation of which is essential for understanding individual carbon emissions and associated health benefits [18; 57].
@@
Traditionally, transportation surveys were used to estimate individual’s choices of transportation mode [63]. However, two major technological advancements in recent years have reshaped transportation mode detection. First, smartphones—equipped with GPS, accelerometers, gyroscopes, and cellular network connectivity—have enabled the creation of detailed mobility datasets derived from mobile applications [41], social media platforms [46], travel cards [21], and cellphone signals [36]. These datasets surpassed traditional transportation surveys in dimensionality, accuracy, variety, and volume [4; 20].
@@
Second, advancements in machine learning (ML) have greatly improved the ability to extract meaningful information from mobility datasets. Machine learning techniques—from ensemble methods to deep learning models—now exploit increasingly fine-grained spatiotemporal data, including human mobility traces, substantially improving the accuracy of transportation-mode prediction [44]. More recently, Transformer [62] architectures are adopted for human mobility research given their self-attention mechanisms, which can help capture nonlinear spatiotemporal dependencies and high-order motion features. Furthermore, their capacity for inductive transfer enables cross-regional generalization, potentially allowing models to adapt to diverse transportation infrastructures using minimal labeled data.
@@
Nevertheless, significant challenges remained. Smartphone-derived mobility datasets, while highly granular, often exhibited inconsistent quality due to the absence of standardized data collection protocols and inaccuracies in geographic information, complicating modeling process. Model performance frequently depended on derived features such as acceleration, while complex preprocessing tasks—such as geocoding and geotagging—further increased data uncertainty. Moreover, extensive feature engineering used to aggregate raw GPS trajectories often resulted in the loss of critical sequential information essential for accurate mobility modeling [29].
@@
Privacy posed another major challenge. Mobility applications commonly collected sensitive information, including geographic locations, trip details (e.g., start and end times), and personal or device identifiers. For example, Xu et al. [65] demonstrated that even anonymized mobility data could be re-identified with 73–91% accuracy, underscoring the difficulty of ensuring privacy protection. Such risks frequently discouraged individuals from sharing their mobility data [49], constraining data availability.
@@
Moreover, mobility models have often lacked generalizability across geographic regions, particularly when trained on isolated or geographically homogeneous datasets. Human mobility behaviors vary widely across countries and regions due to differences in road infrastructure, speed regulations, and cultural norms. Nevertheless, most studies have assessed model performance using train–test splits from the same dataset [72], which fail to capture real-world adaptability. Models fine-tuned on benchmarks such as Geolife [74] frequently required substantial recalibration when applied to other contexts. The proliferation of deep learning frameworks, hyperparameter tuning strategies, and architectural variations has further complicated cross-regional reproducibility in mobility modeling research [22].
@@
Finally, existing mobility models have often struggled to perform reliably under real-world conditions, which are considerably more unpredictable and complex than curated research datasets suggest. Everyday travel involves nuances such as short trips and unexpected errands that traditional models frequently misclassified [44]. Moreover, GPS signal quality varied across smartphone models and was highly sensitive to environmental factors, including urban canyons and signal obstructions [70; 10]. These real-world data inconsistencies—typically underrepresented in benchmark datasets—underscore the need to validate mobility models under realistic conditions that reflect the full variability of human movement.
@@ 2 Related Work
Transportation mode detection represents a central dimension of human mobility research, with downstream applications ranging from carbon footprint estimation [37] and tourism recommendations [64] to traffic management [45] and public health interventions [2]. With the proliferation of GPS-enabled devices and advances in machine learning, the field has undergone substantial progress in recent years, moving beyond traditional social scientific methods such as transportation surveys and adopting data-driven methodologies grounded in machine learning. Notable developments have emerged in both classical machine learning algorithms and deep learning approaches.
@@ 2.1 Machine Learning for Transportation Mode Detection and its Challenges
Classical machine learning (ML) algorithms formed the foundation of early data-driven approaches to transportation mode detection. These methods typically converted sequential GPS trajectories into statistical, tabular representations, with notable implementations including Decision Trees [73], Random Forests [54; 29], and Support Vector Machines [6]. Stenneth et al. [54] demonstrated the effectiveness of these models through systematic evaluation, while Jahangiri and Rakha [29] highlighted their versatility across different transportation modes, variations, and contexts. The logic underlying classical ML algorithms was conceptually similar to that of rule-based models, which relied on indicative features computed from GPS trajectories and used them to infer travel modes statistically. For example, mobility features derived from dense GPS trajectories—such as average speed and acceleration rate—were often effective in distinguishing between driving, walking, and cycling.
@@
Moreover, such methods posed risks to privacy and anonymity, as even a small number of precise and longitudinal GPS points could be sufficient to re-identify individuals [13]. Growing concerns regarding the collection, management, and disclosure of personal GPS data, together with advances in re-identification techniques, have further raised ethical issues surrounding GPS trajectory research [39; 32; 40]. In response, scholars and regulators have increasingly advocated for privacy-preserving techniques in mobility research—including applications such as transportation mode detection—as alternatives to classical machine learning approaches that depend on rich location features and extensive feature engineering [43; 33; 53; 59; 30].
Recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks [25], emerged as a dominant approach for transportation mode prediction due to their capacity to capture temporal dependencies in sequential data. Jiang et al. [31] pioneered the application of LSTMs for mobility analysis, and Asci and Guvensan [3] extended this work by incorporating attention mechanisms that improved performance across varied trip lengths. Hybrid architectures combining LSTMs with other neural components further advanced model accuracy—for example, the Conv-LSTM model proposed by Nawaz et al. [42] leveraged convolutional layers for spatial feature extraction prior to temporal processing. Other innovations included Convolutional Neural Networks (CNNs) [11], autoregressive flow models [17], and semi-supervised learning approaches [12]. Despite these accuracy gains, such sophisticated architectures often required substantial computational resources and complex pre-processing to structure data according to model specifications, thereby limiting their practicality for real-world deployment [22].
@@ 2.2 Transformers for Mobility Modeling
Transformer architectures revolutionized sequence modeling through their reliance on self-attention and multi-head attention mechanisms [9; 62]. Models such as GPT-2 [47] and BERT [14] demonstrated exceptional performance on sequence-based data, significantly surpassing classical machine learning models in capturing complex sequential patterns. This success motivated researchers to explore whether Transformers’ superior capacity for modeling long-range dependencies could similarly advance mobility modeling, where understanding the relationships between distant points within a trajectory is critical [66].
Recent applications of Transformer architectures to mobility studies have shown promise, although the field remains in its early stages. For example, Hong et al. [26] incorporated Transformer components for next-location prediction while treating mode identification as a secondary task. Liang et al. [35] addressed challenges such as irregular spatiotemporal intervals but focused primarily on spatio-temporal dependencies rather than transportation mode prediction. More recently, Ribeiro et al. [48] achieved strong results using a vision Transformer approach [15], although their implementation required converting trajectories into image-based representations. Similarly, Drosouli et al. [16] converted GPS trajectory points into word tokens to apply the original BERT model [14], achieving notable results; however, this approach imposed a linguistic abstraction onto spatial data, making it less suitable for general-purpose mobility modeling.
@@ 3 SpeedTransformer Architecture
We replace the standard sinusoidal positional encoding with Rotary Positional Embeddings (RoPE), which is applied directly to the query ($Q$) and key ($K$) vectors in the attention mechanism [55]. RoPE encodes positional information through position-dependent rotations indexed by the sequence order, enabling the attention mechanism to model relative temporal dependencies in a continuous and rotation-invariant manner. This formulation is particularly well-suited for sequential mobility signals such as speed trajectories, where relative temporal structure is more informative than absolute position.
The encoder consists of $L=4$ Pre-Norm Transformer blocks. Each block contains two key components: a Grouped-Query Attention (GQA) layer [1], which efficiently computes attention by grouping multiple query vectors per key–value pair, and a SwiGLU-activated feed-forward sublayer [52] that introduces non-linear transformations with improved gradient flow and expressivity (see Appendix E for a detailed explanation of SwiGLU). The computation within each block is expressed as follows:
@@ 4 Datasets
We utilized two longitudinal tracking datasets—the widely adopted Geolife dataset [74] and the Swiss MOBIS dataset [24]—which offered complementary strengths for evaluating our model. Summary statistics are presented in Table 1. As illustrated in Figure 2, we standardized their data structures through a unified pre-processing pipeline: we consolidated transportation modes into five consistent categories (Bike, Bus, Car, Train, Walk), unified the trajectory input format across datasets, excluded multi-model trajectories, converted geographic coordinates into speed sequences, removed abnormal trips with erroneous location data, and standardized transportation mode labels. Detailed descriptions of the pre-processing procedures and raw dataset statistics are provided in Appendix A.
@@ 4.1 MOBIS Dataset
The MOBIS dataset [41] was derived from an eight-week randomized controlled trial (RCT) on transport pricing involving 3,680 participants in Switzerland. Each participant used the Catch-my-Day mobile application (available for both iOS and Android), which continuously recorded GPS data via the device’s location services. The application captured daily travel patterns, storing raw trajectory data locally before uploading them to the MotionTag analytics platform, where trip segmentation and transportation mode inference were performed. This extensive data collection process produced 255.3 million GPS records and 1.58 million labeled trips in the raw dataset. After applying our standardized pre-processing pipeline described in Appendix A, the resulting MOBIS dataset contained 144.7 million data points across 1.53 million unique trips distributed among five transportation modes, as summarized in Table 1.
@@ 4.2 Geolife Dataset
Geolife [74] was a widely used benchmark in transportation mode research, collected by Microsoft Research Asia from 182 users in Beijing over a five-year period (2007–2012). The dataset captured urban mobility through 17,000 trajectories covering approximately 1.2 million kilometers across Beijing’s complex transportation network. Data were recorded using various GPS loggers and GPS-enabled phones with different sampling rates, with 91% of trajectories collected at high density (every 1–5 seconds or every 5–10 meters). In addition to routine commutes, the Geolife dataset included leisure and sports activities such as shopping, sightseeing, and cycling, offering rich contextual diversity in trip purposes. After pre-processing to align with the MOBIS data structure, the Geolife dataset contained 4.44 million data points and 9,427 unique trips—substantially smaller than MOBIS, yet providing valuable geographical and temporal diversity for evaluating our model’s performance.
@@ 5 Experiments
We evaluated SpeedTransformer under three experimental conditions. First, we benchmarked it against state-of-the-art transportation mode identification models on Geolife [74], enabling direct comparison with existing approaches. Second, we examined performance consistency across different geographical contexts by comparing it with classical LSTM baseline models on both the Swiss (MOBIS) [24] and Chinese (Geolife) datasets. Finally, we assessed cross-regional transferability through fine-tuning experiments, in which models pretrained on Swiss data were adapted to Chinese mobility patterns using small sample subsets.
@@ 5.1 Benchmarking Performance
  • LSTM-Attention (Baseline): Our reconstructed baseline model implementing a classical bidirectional LSTM with attention mechanism [25]. This baseline is specifically designed to test whether pure attention-based mechanisms outperform recurrent networks augmented with attention, while maintaining the same minimal input requirement (speed only).
  • Deep-ViT [48]: A Vision Transformer approach that transforms GPS features (speed, acceleration, bearing) into image representations using DeepInsight methodology before processing with a Vision Transformer architecture, combining traditional feature engineering with advanced deep learning.
  • CE-RCRF [71]: A sequence-to-sequence framework (TaaS) that processes entire trajectories using a Convolutional Encoder to extract high-level features and a Recurrent Conditional Random Field to maintain contextual information at both feature and label levels, with specialized bus-related features to distinguish high-speed modes.
  • LSTM-based DNN [68]: An ensemble of four LSTM networks that incorporates both time-domain trajectory attributes and frequency-domain statistics developed through discrete Fourier and wavelet transforms, creating a semi-supervised deep learning approach.
  • SECA [12]: A Semi-supervised Convolutional Autoencoder that integrates a convolutional-deconvolutional autoencoder with a CNN classifier to simultaneously leverage labeled and unlabeled GPS segments, automatically extracting relevant features from 4-channel tensor representations.
  • ConvLSTM [42]: A hybrid architecture that uses convolutional layers to extract spatial features from GPS data, followed by LSTM layers to capture temporal patterns, incorporating both location and weather features to enhance mode detection.
@@
Deep-ViT [48] 92.96
LSTM-based DNN [68] 92.70
@@
CE-RCRF [71] 85.23
SECA [12] 84.80
ConvLSTM [42] 83.81

### v2

# Detecting Transportation Mode Using Dense Smartphone GPS Trajectories and Transformer Models

@@ 1 Introduction
The study of human mobility patterns—how individuals move across space—has become an important topic across geography, transportation science, public health, and climate change science [Gonzalez et al., 2008; Schuessler and Axhausen, 2009; Yao et al., 2020; Shaw et al., 2016; Bonaccorsi et al., 2020; McMichael, 2020; Barbosa et al., 2021; Zook et al., 2015; Guo et al., 2015; Tao et al., 2018]. One key aspect of human mobility is transportation mode choice, the accurate estimation of which is essential for understanding individual carbon emissions and associated health benefits [Girod et al., 2013; Tajalli and Hajbabaie, 2017].
@@
Traditionally, transportation surveys were used to estimate individual’s choices of transportation mode [Wang et al., 2015]. However, two major technological advancements in recent years have reshaped transportation mode detection. First, smartphones—equipped with GPS, accelerometers, gyroscopes, and cellular network connectivity—have enabled the creation of detailed mobility datasets derived from mobile applications [Molloy et al., 2023], social media platforms [Preotiuc-Pietro and Cohn, 2013], travel cards [Gordon et al., 2013], and cellphone signals [Lu et al., 2012]. These datasets surpassed traditional transportation surveys in dimensionality, accuracy, variety, and volume [Barbosa et al., 2015; Goodchild, 2013].
@@
Second, advancements in machine learning (ML) have greatly improved the ability to extract meaningful information from mobility datasets. Machine learning techniques—from ensemble methods to deep learning models—now exploit increasingly fine-grained spatiotemporal data, including human mobility traces, substantially improving the accuracy of transportation-mode prediction [Pappalardo et al., 2023]. More recently, Transformer [Vaswani et al., 2017] architectures are adopted for human mobility research given their self-attention mechanisms, which can help capture nonlinear spatiotemporal dependencies and high-order motion features. Furthermore, their capacity for inductive transfer enables cross-regional generalization, potentially allowing models to adapt to diverse transportation infrastructures using minimal labeled data.
@@
Nevertheless, significant challenges remained. Smartphone-derived mobility datasets, while highly granular, often exhibited inconsistent quality due to the absence of standardized data collection protocols and inaccuracies in geographic information, complicating modeling process. Model performance frequently depended on derived features such as acceleration, while complex preprocessing tasks—such as geocoding and geotagging—further increased data uncertainty. Moreover, extensive feature engineering used to aggregate raw GPS trajectories often resulted in the loss of critical sequential information essential for accurate mobility modeling [Jahangiri and Rakha, 2015].
@@
Privacy posed another major challenge. Mobility applications commonly collected sensitive information, including geographic locations, trip details (e.g., start and end times), and personal or device identifiers. For example, Xu et al. [2017] demonstrated that even anonymized mobility data could be re-identified with 73–91% accuracy, underscoring the difficulty of ensuring privacy protection. Such risks frequently discouraged individuals from sharing their mobility data [Rzeszewski and Luczys, 2018], constraining data availability.
@@
Moreover, mobility models have often lacked generalizability across geographic regions, particularly when trained on isolated or geographically homogeneous datasets. Human mobility behaviors vary widely across countries and regions due to differences in road infrastructure, speed regulations, and cultural norms. Nevertheless, most studies have assessed model performance using train–test splits from the same dataset [Zhao et al., 2016], which fail to capture real-world adaptability. Models fine-tuned on benchmarks such as Geolife [Zheng et al., 2010] frequently required substantial recalibration when applied to other contexts. The proliferation of deep learning frameworks, hyperparameter tuning strategies, and architectural variations has further complicated cross-regional reproducibility in mobility modeling research [Graser et al., 2024].
@@
Finally, existing mobility models have often struggled to perform reliably under real-world conditions, which are considerably more unpredictable and complex than curated research datasets suggest. Everyday travel involves nuances such as short trips and unexpected errands that traditional models frequently misclassified [Pappalardo et al., 2023]. Moreover, GPS signal quality varied across smartphone models and was highly sensitive to environmental factors, including urban canyons and signal obstructions [Zandbergen, 2009; Cui and Ge, 2001]. These real-world data inconsistencies—typically underrepresented in benchmark datasets—underscore the need to validate mobility models under realistic conditions that reflect the full variability of human movement.
@@ 2 Related Work
Transportation mode detection represents a central dimension of human mobility research, with downstream applications ranging from carbon footprint estimation [Manzoni et al., 2010] and tourism recommendations [Xiao et al., 2025] to traffic management [Prelipcean et al., 2016] and public health interventions [Aleta et al., 2022]. With the proliferation of GPS-enabled devices and advances in machine learning, the field has undergone substantial progress in recent years, moving beyond traditional social scientific methods such as transportation surveys and adopting data-driven methodologies grounded in machine learning. Notable developments have emerged in both classical machine learning algorithms and deep learning approaches.
@@ 2.1 Machine Learning for Transportation Mode Detection and its Challenges
Classical machine learning (ML) algorithms formed the foundation of early data-driven approaches to transportation mode detection. These methods typically converted sequential GPS trajectories into statistical, tabular representations, with notable implementations including Decision Trees [Zheng et al., 2008], Random Forests [Stenneth et al., 2011; Jahangiri and Rakha, 2015], and Support Vector Machines [Bolbol et al., 2012a]. Stenneth et al. [2011] demonstrated the effectiveness of these models through systematic evaluation, while Jahangiri and Rakha [2015] highlighted their versatility across different transportation modes, variations, and contexts. The logic underlying classical ML algorithms was conceptually similar to that of rule-based models, which relied on indicative features computed from GPS trajectories and used them to infer travel modes statistically. For example, mobility features derived from dense GPS trajectories—such as average speed and acceleration rate—were often effective in distinguishing between driving, walking, and cycling.
@@
Moreover, such methods posed risks to privacy and anonymity, as even a small number of precise and longitudinal GPS points could be sufficient to re-identify individuals [De Montjoye et al., 2013]. Growing concerns regarding the collection, management, and disclosure of personal GPS data, together with advances in re-identification techniques, have further raised ethical issues surrounding GPS trajectory research [Michael et al., 2006; Klasnja et al., 2009; Minch, 2004]. In response, scholars and regulators have increasingly advocated for privacy-preserving techniques in mobility research—including applications such as transportation mode detection—as alternatives to classical machine learning approaches that depend on rich location features and extensive feature engineering [Ng-Kruelle et al., 2002; Krumm, 2009; Shin et al., 2012; Thompson and Warzel, 2022; Jiang et al., 2021].
Recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks [Hochreiter and Schmidhuber, 1997], emerged as a dominant approach for transportation mode prediction due to their capacity to capture temporal dependencies in sequential data. Jiang et al. [2017] pioneered the application of LSTMs for mobility analysis, and Asci and Guvensan [2019] extended this work by incorporating attention mechanisms that improved performance across varied trip lengths. Hybrid architectures combining LSTMs with other neural components further advanced model accuracy—for example, the Conv-LSTM model proposed by Nawaz et al. [2020] leveraged convolutional layers for spatial feature extraction prior to temporal processing. Other innovations included Convolutional Neural Networks (CNNs) [Dabiri and Heaslip, 2018], autoregressive flow models [Dutta and Patra, 2023], and semi-supervised learning approaches [Dabiri et al., 2020]. Despite these accuracy gains, such sophisticated architectures often required substantial computational resources and complex pre-processing to structure data according to model specifications, thereby limiting their practicality for real-world deployment [Graser et al., 2024].
@@ 2.2 Transformers for Mobility Modeling
Transformer architectures revolutionized sequence modeling through their reliance on self-attention and multi-head attention mechanisms [Cheng et al., 2016; Vaswani et al., 2017]. Models such as GPT-2 [Radford et al., 2019] and BERT [Devlin et al., 2019] demonstrated exceptional performance on sequence-based data, significantly surpassing classical machine learning models in capturing complex sequential patterns. This success motivated researchers to explore whether Transformers’ superior capacity for modeling long-range dependencies could similarly advance mobility modeling, where understanding the relationships between distant points within a trajectory is critical [Xue et al., 2022].
Recent applications of Transformer architectures to mobility studies have shown promise, although the field remains in its early stages. For example, Hong et al. [2022] incorporated Transformer components for next-location prediction while treating mode identification as a secondary task. Liang et al. [2022] addressed challenges such as irregular spatiotemporal intervals but focused primarily on spatio-temporal dependencies rather than transportation mode prediction. More recently, Ribeiro et al. [2024] achieved strong results using a vision Transformer approach [Dosovitskiy et al., 2021], although their implementation required converting trajectories into image-based representations. Similarly, Drosouli et al. [2023] converted GPS trajectory points into word tokens to apply the original BERT model [Devlin et al., 2019], achieving notable results; however, this approach imposed a linguistic abstraction onto spatial data, making it less suitable for general-purpose mobility modeling.
@@ 3 SpeedTransformer Architecture
We replace the standard sinusoidal positional encoding with Rotary Positional Embeddings (RoPE), which is applied directly to the query ($Q$) and key ($K$) vectors in the attention mechanism [Su et al., 2021]. RoPE encodes positional information through position-dependent rotations indexed by the sequence order, enabling the attention mechanism to model relative temporal dependencies in a continuous and rotation-invariant manner. This formulation is particularly well-suited for sequential mobility signals such as speed trajectories, where relative temporal structure is more informative than absolute position.
The encoder consists of $L=4$ Pre-Norm Transformer blocks. Each block contains two key components: a Grouped-Query Attention (GQA) layer [Ainslie et al., 2023], which efficiently computes attention by grouping multiple query vectors per key–value pair, and a SwiGLU-activated feed-forward sublayer [Shazeer, 2020] that introduces non-linear transformations with improved gradient flow and expressivity (see Appendix E for a detailed explanation of SwiGLU). The computation within each block is expressed as follows:
@@ 4 Datasets
We utilized two longitudinal tracking datasets—the widely adopted Geolife dataset [Zheng et al., 2010] and the Swiss MOBIS dataset [Heimgartner and Axhausen, 2024]—which offered complementary strengths for evaluating our model. Summary statistics are presented in Table 1. As illustrated in Figure 2, we standardized their data structures through a unified pre-processing pipeline: we consolidated transportation modes into five consistent categories (Bike, Bus, Car, Train, Walk), unified the trajectory input format across datasets, excluded multi-model trajectories, converted geographic coordinates into speed sequences, removed abnormal trips with erroneous location data, and standardized transportation mode labels. Detailed descriptions of the pre-processing procedures and raw dataset statistics are provided in Appendix A.
@@ 4.1 MOBIS Dataset
The MOBIS dataset [Molloy et al., 2023] was derived from an eight-week randomized controlled trial (RCT) on transport pricing involving 3,680 participants in Switzerland. Each participant used the Catch-my-Day mobile application (available for both iOS and Android), which continuously recorded GPS data via the device’s location services. The application captured daily travel patterns, storing raw trajectory data locally before uploading them to the MotionTag analytics platform, where trip segmentation and transportation mode inference were performed. This extensive data collection process produced 255.3 million GPS records and 1.58 million labeled trips in the raw dataset. After applying our standardized pre-processing pipeline described in Appendix A, the resulting MOBIS dataset contained 144.7 million data points across 1.53 million unique trips distributed among five transportation modes, as summarized in Table 1.
@@ 4.2 Geolife Dataset
Geolife [Zheng et al., 2010] was a widely used benchmark in transportation mode research, collected by Microsoft Research Asia from 182 users in Beijing over a five-year period (2007–2012). The dataset captured urban mobility through 17,000 trajectories covering approximately 1.2 million kilometers across Beijing’s complex transportation network. Data were recorded using various GPS loggers and GPS-enabled phones with different sampling rates, with 91% of trajectories collected at high density (every 1–5 seconds or every 5–10 meters). In addition to routine commutes, the Geolife dataset included leisure and sports activities such as shopping, sightseeing, and cycling, offering rich contextual diversity in trip purposes. After pre-processing to align with the MOBIS data structure, the Geolife dataset contained 4.44 million data points and 9,427 unique trips—substantially smaller than MOBIS, yet providing valuable geographical and temporal diversity for evaluating our model’s performance.
@@ 5 Experiments
We evaluated SpeedTransformer under three experimental conditions. First, we benchmarked it against state-of-the-art transportation mode identification models on Geolife [Zheng et al., 2010], enabling direct comparison with existing approaches. Second, we examined performance consistency across different geographical contexts by comparing it with classical LSTM baseline models on both the Swiss (MOBIS) [Heimgartner and Axhausen, 2024] and Chinese (Geolife) datasets. Finally, we assessed cross-regional transferability through fine-tuning experiments, in which models pretrained on Swiss data were adapted to Chinese mobility patterns using small sample subsets.
@@ 5.1 Benchmarking Performance
  • LSTM-Attention (Baseline): Our reconstructed baseline model implementing a classical bidirectional LSTM with attention mechanism [Hochreiter and Schmidhuber, 1997]. This baseline is specifically designed to test whether pure attention-based mechanisms outperform recurrent networks augmented with attention, while maintaining the same minimal input requirement (speed only).
  • Deep-ViT [Ribeiro et al., 2024]: A Vision Transformer approach that transforms GPS features (speed, acceleration, bearing) into image representations using DeepInsight methodology before processing with a Vision Transformer architecture, combining traditional feature engineering with advanced deep learning.
  • CE-RCRF [Zeng et al., 2023]: A sequence-to-sequence framework (TaaS) that processes entire trajectories using a Convolutional Encoder to extract high-level features and a Recurrent Conditional Random Field to maintain contextual information at both feature and label levels, with specialized bus-related features to distinguish high-speed modes.
  • LSTM-based DNN [Yu, 2020]: An ensemble of four LSTM networks that incorporates both time-domain trajectory attributes and frequency-domain statistics developed through discrete Fourier and wavelet transforms, creating a semi-supervised deep learning approach.
  • SECA [Dabiri et al., 2020]: A Semi-supervised Convolutional Autoencoder that integrates a convolutional-deconvolutional autoencoder with a CNN classifier to simultaneously leverage labeled and unlabeled GPS segments, automatically extracting relevant features from 4-channel tensor representations.
  • ConvLSTM [Nawaz et al., 2020]: A hybrid architecture that uses convolutional layers to extract spatial features from GPS data, followed by LSTM layers to capture temporal patterns, incorporating both location and weather features to enhance mode detection.
@@
Deep-ViT [Ribeiro et al., 2024] 92.96
LSTM-based DNN [Yu, 2020] 92.70
@@
CE-RCRF [Zeng et al., 2023] 85.23
SECA [Dabiri et al., 2020] 84.80
ConvLSTM [Nawaz et al., 2020] 83.81